OpenStudy (butterflydreamer):

Solve the following first order difference equation:

OpenStudy (butterflydreamer):

\[2x_{n+1} + 3x_{n} = 0 \]

OpenStudy (kainui):

Are you expected to use certain methods, or just do whatever works?

OpenStudy (butterflydreamer):

I don't really understand how to solve these type of problems. I know that the solution is meant to be something along the lines of \[x_{n} = C(a^n)\]

OpenStudy (kainui):

You can algebraically rearrange it to be: \[x_{n+1} = \left(\frac{-3}{2}\right) x_n\] So if you think of it as saying "the next term in the sequence is a multiple of the previous term" it sorta makes sense that it would be of the form: \[x_n = C*a^n\] where C is some arbitrary constant and 'a' is some number. Each time you increase the exponent n by 1, it multiplies by a. C is like some base starting amount and I think you'll need more information to determine it.

OpenStudy (butterflydreamer):

ahh okay. I think i'm starting to kind of understand it. Thanks :)

OpenStudy (kainui):

There are other ways of doing this that are more like the strategy a computer could use, these would be like generating functions. There's really not a simple way of doing these in a way that "oh ok this strategy works for all of them all the time" kinda thing unfortunately. I wouldn't have been able to tell you the answer so quickly if I hadn't had someone else tell me this kinda stuff.

OpenStudy (butterflydreamer):

Wait so if a difference equation is in the form : \[x_n - ax_{n-1} = 0 \] where a is a constant, the solution will be in the form: \[x_n = Ca^n\] I think i saw somewhere.. that a is the "root of the auxiliary equation m-a = 0" So in saying so, Does that mean we write something like; \[x_{n+1} \rightarrow m\] \[x_n \rightarrow -1\] \[2m - 3 = 0\] \[m = \frac{ -3 }{ 2 }\] so then \[a = \frac{ -3 }{ 2 }\] and then we sub that into x_n = Ca^n so the solution would be: \[x_n = C(\frac{ -3 }{ 2 })^n\] ???

OpenStudy (kainui):

I am not sure about this 'auxiliary equation' terminology I'm just not familiar with it. However you can always check a solution: I just like the forward difference better than backwards difference but it really makes no difference what you choose... So if you want to solve \(x_{n+1} = a x_n\) you can check a guess: \[x_n=Ca^n\] by plugging it in: \[x_{n+1} = Ca^{n+1}\] \[x_{n+1} = a x_n\]\[Ca^{n+1} = a Ca^n\] Yeah it checks out. If you do the same thing with your specific choice of solution, you'll see that it indeed works. This method is probably a good thing to do after every problem you solve just in case. It's a simple check to plug in your solution.

OpenStudy (butterflydreamer):

okay cool. Thanks for your help! :)